OpenStudy (anonymous):

solve the system by the substitution method 1/2x+1/3y= -1/2 1/2x+3y=15/2 the / make them fractions

OpenStudy (anonymous):

Solve for y, then substitute the equations y = y, then solve for x

OpenStudy (anonymous):

i would multiply first and second equation by 6 to get rid of annoying fractions first

OpenStudy (anonymous):

multiply the two equations by six first of all

OpenStudy (anonymous):

actually i lied

OpenStudy (anonymous):

i would multiply first by 6, second by 2

OpenStudy (anonymous):

get \[3x+2y=-3\] \[x+6y=15\]

OpenStudy (anonymous):

then if you really want "substitution" write second one as \[x=-6y+15\] and substitute into first equation to get \[3(-6y+15)+2y=-3\]

OpenStudy (anonymous):

\[-18y+45+2y=-3\] \[-16y=-48\] \[y=\frac{48}{16}=3\]

OpenStudy (anonymous):

I guess nobody liked my method

OpenStudy (anonymous):

then substitute back to get \[x+6\times 3=15\][ \[x=18=15\] \[x=-3\] and so your answer is (-3,3)

OpenStudy (anonymous):

no i didn't like your method because fraction annoy me. but of course it makes no difference.

OpenStudy (anonymous):

It's not a big deal. Just get it in terms of y, set the equations equal to each other, then cross multiply. It's actually less steps

OpenStudy (anonymous):

ok write it out and see which one is easier to write